OpenStudy (anonymous):

What is the distance between -4,0 and 5,0?

OpenStudy (anonymous):

\[\sqrt{(y _{2}-y _{1})^2-(x _{2}-x _{1})^2}\]

OpenStudy (anonymous):

use this

OpenStudy (owlcoffee):

Distance formula between two poins P1(x1,y1) P2(x2,y2): \[d(p_1p_2)=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

*+

OpenStudy (anonymous):

im extermely sorry :)

OpenStudy (anonymous):

it would substitute in as (5+4)^2 + (0-0)^2

OpenStudy (owlcoffee):

\[d(p_1p_2)=\sqrt{(9)^2}\]

OpenStudy (anonymous):

the reason why it is (5+4) is because 2 negatives make a positive

OpenStudy (owlcoffee):

example: \[b-(-a)=b+a\]

OpenStudy (anonymous):

then (5+4)^2 would be 9^2 and (0-0)^2 would be 0^2

OpenStudy (anonymous):

9^2 would turn into 81 and 0^2 would turn into 0

OpenStudy (anonymous):

81+0=81

OpenStudy (anonymous):

then you will have to find the square root of 81

OpenStudy (owlcoffee):

\[d(p_1p_2)=\sqrt{81}\]

OpenStudy (anonymous):

which is 9

OpenStudy (anonymous):

so your answer will be 9

OpenStudy (owlcoffee):

\[d(p_1p_2)=9\]

OpenStudy (anonymous):

yes correct :) Good luck!!